OpenStudy (anonymous):

Find the constant c so that a random variable X is defined by taking on values 1 , 2 , … , n , … with probabilities c*(3/5)^n

OpenStudy (anonymous):

I was thinking \[\frac{3}{5} = 1\] for the constant so \[1-\frac{3}{5} = \frac{2}{5}\] \[\frac{\frac{3}{5}}{\frac{2}{5}} = \frac{3}{2}\]

OpenStudy (anonymous):

is \[\frac{3}{2}= 1\] \[c=\frac{2}{3}\]?

OpenStudy (anonymous):

is my solution correct ?